OpenStudy (anonymous):

Faith observed 16 bugs while doing research for her science project. She observed insects and spiders, which had a total of 114 legs. Bearing in mind that insects have 6 legs and spiders have 8 legs, how many insects did she observe? A. 7 B. 8 C. 9 D. 10

hero (hero):

x = spiders y = insects x + y = 16 8x + 6y = 114

hero (hero):

Can you solve from there or no?

OpenStudy (anonymous):

What numbers could I use to substitue for x and y?

OpenStudy (anonymous):

7, 8, 9, 10?

hero (hero):

y = 16 - x y = (114-8x)/6 y = y 6(16 - x) = 114 - 8x 96 - 6x = 114 - 8x 8x - 6x = 114 - 96 2x = 18 x = 9 y = 7

OpenStudy (istim):

I think you should try them all?

hero (hero):

so 7 insects total

hero (hero):

Well, amongst them, you know that they should add to get 16 so if you had to guess, it would be between 7 and 9 as far as guessing answer choices

hero (hero):

But using algebra, we don't have to guess

OpenStudy (anonymous):

Lol okay. Thank you Hero :-)